# Utilizing and extending superconducting circuit toolbox to simulate analog quantum gravity

### v1

@@ Appendix D Correlation matrix from Klich’s determinant formula
$=e^{-\text{tr}A^{(0)}}\det\left(\frac{\mathbb{I}_{2J}+Z}{2}+Q^{-1}e^{ZA}Q\frac{\mathbb{I}_{2J}-Z}{2}\right)^{-1}$ (145)
@@
$=$ $\lim_{\delta\chi\to 0}\frac{\sqrt{\lim_{\beta\to\infty}\frac{\text{Tr}\left(e^{i\mathcal{H}_{>}t}e^{\hat{\chi}}e^{-i\mathcal{H}_{>}t}e^{-\beta\mathcal{H}_{<}}\right)^{2}}{\text{Tr}\left(e^{-\beta\mathcal{H}_{<}}\right)^{2}}}-1}{\delta\chi}$ (146)
@@
$\hat{\chi}$ $=\delta\chi\left(\frac{1}{2}\mathbf{a}^{\dagger}M\mathbf{a}\pm\frac{1}{2}\text{tr}M^{(0)}\right)$
$M$ $=\left(\begin{array}[]{cc}M^{(0)}&M^{(1)}\\ M^{(2)}&\left(M^{(0)}\right)^{T}\end{array}\right)$ (149)
@@
$M^{(1)}=$ $\left(M^{(1)}\right)^{T}$
$M^{(2)}=$ $\left(M^{(2)}\right)^{T}$
@@
$=\lim_{\delta\chi\to 0}\frac{\sqrt{\lim_{\beta\to\infty}\frac{e^{\pm\delta\chi\text{tr}M^{(0)}}\det\left(1-e^{iZ_{2J}H_{>}t}e^{\delta\chi Z_{2J}M}e^{-iZ_{2J}H_{>}t}e^{-\beta Z_{2J}H_{<}}\right)^{-1}}{\det\left(1-e^{-\beta Z_{2J}H_{<}}\right)^{-1}}}-1}{\delta\chi}$ (150)
@@
    $M_{lk}^{(0)}$ =$=\begin{cases}0&l\neq i,k\neq j\\ 1&l=i,k=j\end{cases}$
    $\hat{\chi}$ $=\delta\chi\left(\frac{1}{2}\mathbf{a}^{\dagger}M\mathbf{a}-\frac{1}{2}\text{tr}M^{(0)}\right)$ (156)
@@
    $M_{lk}^{(1)}$ =$=\begin{cases}1&l=j-J,k=i\\ 1&l=i,k=j-J\\ 0&\text{every other element}\end{cases}$
@@
    $M_{lk}^{(2)}$ =$=\begin{cases}1&l=j,k=i-J\\ 1&l=i-J,k=j\\ 0&\text{every other element}\end{cases}$
@@
    $M_{lk}^{(0)}$ =$=\begin{cases}0&l\neq i-J,k\neq j-J\\ 1&l=i-J,k=j-J\end{cases}$
    $\hat{\chi}$ $=\delta\chi\left(\frac{1}{2}\mathbf{a}^{\dagger}M\mathbf{a}+\frac{1}{2}\text{tr}M^{(0)}\right)$ (159)

### v2

@@ Appendix D Correlation matrix from Klich’s determinant formula
$=e^{-\text{tr}A^{(0)}}\det\left(\frac{\mathbb{I}_{2J}+Z}{2}+Q^{-1}e^{ZA}Q\frac{\mathbb{I}_{2J}-Z}{2}\right)^{-1},$ (145)
@@
$=$ $\lim_{\delta\chi\to 0}\frac{\sqrt{\lim_{\beta\to\infty}\frac{\text{Tr}\left(e^{i\mathcal{H}_{>}t}e^{\hat{\chi}}e^{-i\mathcal{H}_{>}t}e^{-\beta\mathcal{H}_{<}}\right)^{2}}{\text{Tr}\left(e^{-\beta\mathcal{H}_{<}}\right)^{2}}}-1}{\delta\chi},$ (146)
@@
$\hat{\chi}$ $=\delta\chi\left(\frac{1}{2}\mathbf{a}^{\dagger}M\mathbf{a}\pm\frac{1}{2}\text{tr}M^{(0)}\right),$
$M$ $=\left(\begin{array}[]{cc}M^{(0)}&M^{(1)}\\ M^{(2)}&\left(M^{(0)}\right)^{T}\end{array}\right),$ (149)
@@
$M^{(1)}=$ $\left(M^{(1)}\right)^{T},$
$M^{(2)}=$ $\left(M^{(2)}\right)^{T}.$
@@
$=\lim_{\delta\chi\to 0}\frac{\sqrt{\lim_{\beta\to\infty}\frac{e^{\pm\delta\chi\text{tr}M^{(0)}}\det\left(1-e^{iZ_{2J}H_{>}t}e^{\delta\chi Z_{2J}M}e^{-iZ_{2J}H_{>}t}e^{-\beta Z_{2J}H_{<}}\right)^{-1}}{\det\left(1-e^{-\beta Z_{2J}H_{<}}\right)^{-1}}}-1}{\delta\chi},$ (150)
@@
    $M_{lk}^{(0)}$ =$=\begin{cases}0&l\neq i,k\neq j,\\ 1&l=i,k=j,\end{cases}$
    $\hat{\chi}$ $=\delta\chi\left(\frac{1}{2}\mathbf{a}^{\dagger}M\mathbf{a}-\frac{1}{2}\text{tr}M^{(0)}\right),$ (156)
@@
    $M_{lk}^{(1)}$ =$=\begin{cases}1&l=j-J,k=i,\\ 1&l=i,k=j-J,\\ 0&\text{every other element,}\end{cases}$
@@
    $M_{lk}^{(2)}$ =$=\begin{cases}1&l=j,k=i-J,\\ 1&l=i-J,k=j,\\ 0&\text{every other element,}\end{cases}$
@@
    $M_{lk}^{(0)}$ =$=\begin{cases}0&l\neq i-J,k\neq j-J,\\ 1&l=i-J,k=j-J,\end{cases}$
    $\hat{\chi}$ $=\delta\chi\left(\frac{1}{2}\mathbf{a}^{\dagger}M\mathbf{a}+\frac{1}{2}\text{tr}M^{(0)}\right).$ (159)